Fermi National Laboratory


Accelerator Update for March 4, 2005 - March 7, 2005

Friday March 4
The day shift began with Operations monitoring store 4021 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

Operations placed a single NuMI event in the timeline at 8:46 AM to check out the beamline before the experiment's dedication this afternoon.

The Meson MTest experiment asked for beam at 9:50 AM. Operators had trouble establishing beam to the experiment and so asked an expert for help.

Tevatron (TeV) experts began an end-of-store study at 3:59 PM.

The MTest experiment reported 4:41 PM that they were done with beam for the weekend.

Operations terminated store 4021 at 5:05 PM.

At 5:37 PM, Operations reported having a problem putting the TeV into shot setup due to a Lambertson power supply (LAM52).

Experts began investigating the LAM52 problem at 6:12 PM. They replaced a timing card and had the device fixed by 6:38 PM. Shot setup began.

Operations established store 4022 at 8:35 PM with an initial luminosity of 93.83E30. Pbar resumed stacking.

Operations resumed sending beam to MCenter at 9:43 PM.

At 9:52 PM, Operations reported that the Pbar stacking production was down. Operators investigated.

Saturday March 5
The midnight shift began with Operations monitoring store 4022, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations put the Recycler into shot setup to stash antiprotons at 2:41 AM. Any shot setup halts beam to Switchyard (SY) and NuMI.

Pbar resumed stacking at 3:42 AM.

MCenter resumed taking beam 3:52 AM.

At 8:19 AM, an Accumulator sextupole power supply (A:SEX12) tripped off. This caused the loss of 40mA of antiprotons. Operators called in an expert.

At 8:45 AM, Operations reported that a Linac quadrupole power supply (L:QPS113) had a wandering current. A minute later, QPS114 showed similar problems. Operators investigated and discovered a tripped breaker.

An expert repaired the phase imbalance circuitry for A:SEX12 at 9:35 AM.

Pbar resumed stacking at 9:42 AM.

At 12:37 PM a vertical correction power supply (I:V701) in the P1 Main Injector (MI) transfer line tripped off causing MI to loose its beam permit. It reset okay.

At 1:21 PM, I:V701 tripped again. Operations called in an expert who investigated, reset one of the two power supplies, at 2:38 PM, used to power the string of V701 magnets. Pbar resumed stacking.

NuMI studiers finished using beam around 10 PM.

Operations terminated store 4022 at 10:04 PM and then put the TeV into shot setup a few minutes later.

Sunday March 6
The midnight shift began with the TeV in shot setup, with SY beam off due to shot setup, and with MiniBooNE taking beam.

Operations established store 4023 at 12:46 AM with an initial luminosity of 104.72E30. Pbar resumed stacking.

MCenter resumed taking beam at 12:53 AM.

At 2:33 AM, the MI beam permit tripped off due to a MECAR trip. It reset okay.

Operations put the Recycler into shot setup at 8:11 AM. Antiproton transfers were completed by 9:10 AM. Pbar resumed stacking and MCenter resumed taking beam.

NuMI experts began a study period at 9:59 AM.

At 3:49 PM, the experimenters at D-Zero (D0) reported that their solenoid had tripped off due to a quench.

At 4:34 PM, a Booster kicker (B:MKS02) tripped off. Operators investigated and discovered the area to be very warm. They cooled the area and had the kicker back on by 5:10 PM.

At 5:23 PM, a Booster RF station (BRF4) tripped off. It reset okay.

At 5:37 PM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum pump trip. Everything reset okay.

At 6:14 PM, BRF15 tripped off. It reset okay.

At 7:57 PM, the TeV suffered a sector C4 quench due to a quadrupole power supply (C:D0Q3) trip.

Pbar experts began study around 8 PM.

Operators made D0 ready for an access at 9:17 PM.

NuMI studies said that they were finished with the beam for the night at 9:20 PM.

At 10:07 PM, another Booster kicker (MKS05) tripped off due to high temperature. Operators cooled the area and had the kicker on about ten minutes later.

D0 experimenters came out of access at 11:44 PM.

Monday March 7
The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations put the TeV through a dry squeeze at 12:14 AM.

Operations put the TeV into shot setup at 1:31 AM.

Operations established store 4024 at 4:20 AM with an initial luminosity of 98.07E30. Pbar resumed stacking.

MCenter resumed taking beam at 4:31 AM.

The Plan for Monday
The plan for today is to allow Pbar experts to conduct an antiproton study. The present store will remain colliding until Tuesday (3/8/05) afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/7/2005   email Fermilab